# Implicit Third-Order Peer Triplets with Variable Stepsizes for Gradient-Based Solutions in Large-Scale ODE-Constrained Optimal Control

### v1

@@ Abstract
This paper is concerned with the theory, construction and application of variable-stepsize implicit Peer two-step methods that are super-convergent for variable stepsizes, i.e., preserve their classical order achieved for uniform stepsizes when applied in a gradient-based solution algorithm to solve ODE-constrained optimal control problems in a first-discretize-then-optimize setting. Gradients of the objective function can be computed most efficiently using approximate adjoint variables. High accuracy with moderate computational effort can be achieved through time integration methods that satisfy a sufficiently large number of adjoint order conditions for variable stepsizes and provide gradients with higher-order consistency. In this paper, we enhance our previously developed variable implicit two-step Peer triplets constructed in [J. Comput. Appl. Math. 460, 2025] to get ready for large-scale dynamical systems with varying time scales without loosing efficiency. A key advantage of Peer methods is their use of multiple stages with the same high stage order, which prevents order reduction — an issue commonly encountered in semi-discretized PDE problems with boundary control. Two third-order methods with four stages, good stability properties, small error constants, and a grid adaptation by equi-distributing global errors are constructed and tested for a 1D boundary heat control problem and an optimal control of cytotoxic therapies in the treatment of prostate cancer.
@@ 1 Introduction
The paper is organized as follows. In Section 2, we introduce a generalized form of a Peer two-step triplet. The continuous and discrete first order optimality conditions are presented in Section 3. The framework of gradient-based optimization, which relies on an iterative forward-backward marching scheme, is discussed in Section 4. Several properties of Peer triplets, including the standard and boundary methods, are described in Section 5. In Section 6, we modify two of our previous Peer two-step triplets of practical interest. Grid improvement is discussed in Section  7 and numerical results are given in Section 8. Conclusions are stated in Section 9.
Throughout the paper, we use the following notations: $e_{i}$ denotes a cardinal basis vector, $\mbox{1\thinspace l}_{k}=(1,\ldots,1)^{\sf T}\in{\mathbb{R}}^{k}$ is the vector of ones, and $I_{k}\in{\mathbb{R}}^{k\times k}$ is the identity matrix. The latter two are occasionally used without index if there is no ambiguity.
@@ 3 First order optimality conditions
Here, $p_{h}(T)=\nabla_{y}{{\cal C}}\big{(}y_{h}(T)\big{)}^{\sf T}$ and the Jacobians of $F$ are block diagonal matrices $\nabla_{Y}F(Y_{n},U_{n})=\,\mbox{blockdiag}_{i}\big{(}\nabla_{y}f(Y_{ni},U_{ni})\big{)}$. Since $h_{n}\kappa_{ii}>0$, (15) can be simplified to $\nabla_{u}f(Y_{ni},U_{ni})^{\sf T}P_{ni}\in N_{U}(U_{ni})$. Hence, the control uniqueness property mentioned above guarantees the existence of a local minimizer $U_{ni}$ of the Hamiltonian $H(Y_{ni},U,P_{ni})$ over all $U\in U_{ad}$, which shows the importance of having $\kappa_{ii}>0$ for all diagonal entries of $K$. Such positivity conditions also arise in the context of classical Runge-Kutta methods or W-methods, see [9, Theorem 2.1] and [20, Chapter 5.2].
@@ 5.1 The standard method
As for multi-step methods, order conditions are derived by Taylor expansion of the residuals if exact solution values are used as arguments of the numerical scheme. For variable stepsizes, these conditions depend on the stepsize ratio $\sigma_{n}=h_{n}/h_{n-1}$ in the scaling matrix $S_{n,q}=\,\mbox{diag}_{i}\big{(}\sigma_{n}^{i-1}\big{)}\in{\mathbb{R}}^{q\times q}$. Compact formulations of the order conditions are possible with the vector of node powers ${\bf c}^{k}=(c_{1}^{k},\ldots,c_{s}^{k})^{\sf T}$, the Vandermonde matrix $V_{q}=(\mbox{1\thinspace l},{\bf c},\ldots,{\bf c}^{q-1})\in{\mathbb{R}}^{s\times q}$, the Pascal matrix ${\cal P}_{q}=\big{(}{j-1\choose i-1}\big{)}\in{\mathbb{R}}^{q\times q}$, and the scaled shift matrix $\tilde{E}_{1}=\big{(}i\delta_{i+1,j}\big{)}\in{\mathbb{R}}^{q\times q}$ which commutes with ${\cal P}_{q}=\exp(\tilde{E}_{q})$. Both the standard forward step (12) and its adjoint (14) posses local order $q$ if
@@
A critical requirement for the standard method is zero stability which may be very demanding for variable stepsizes. For the trivial ODE $y^{\prime}=0$, the stability matrix of the forward step (12), $Y_{n}=A^{-1}B_{n}Y_{n-1}$, of the standard method is $\bar{B}_{n}:=A^{-1}B_{n}$ and for the adjoint step (14) it is $A^{\sf-T}B_{n+1}^{\sf T}=A^{\sf-T}\bar{B}_{n+1}^{\sf T}A^{\sf T}$. Zero stability requires that arbitrarily long products, e.g., $\bar{B}_{n}\cdots\bar{B}_{k+1},\,n>k,$ are uniformly bounded. Since $\hat{A}^{-1}\hat{B}(\sigma)=V_{s}^{-1}\bar{B}(\sigma)V_{s}$ is similar to the stability matrix $\bar{B}(\sigma)$, based on the representation (22) in [19], the construction of a $\sigma$-independent weight matrix $W\in{\mathbb{R}}^{s\times s}$ was incorporated into the design of all standard methods, which block-diagonalizes all matrices simultaneously,
@@
$\tau_{n}^{Y}=$ $\;h_{n}^{q}\beta_{q,n}(\sigma_{n})y^{\star(q)}(t_{n})+O\left(h_{n}^{q+1}\|y^{\star(q+1)}\|_{[n]}\right),$ (28)
@@
$\tau_{n}^{P}=$ $\;h_{n}^{q}\beta_{q,n}^{\dagger}(\sigma_{n+1})p^{\star(q)}(t_{n})+O\left(h_{n}^{q+1}\|p^{\star(q+1)}\|_{[n]}\right),$ (30)
@@
see (23), with an analogous version for the adjoint matrices $A^{\sf T}B_{n+1}$. Hence, there is no accumulation of leading error terms of the standard method under the super-convergence conditions
@@ 6.2 The triplet AP4o33vsi for smooth grids
The diagonal matrix $K$ is also positive definite and, hence, preserves the sign in the cone condition (15). The triplet has a substantially larger angle $\alpha>83^{o}$ of $A(\alpha)$-stability than AP4o33vgi and the complement of its stability region is contained in $0.6,20]\times[-15,15]$. One drawback is larger error constants, see Table 1. The standard method is uniformly zero stable in $\underline{\sigma},\bar{\sigma}]=[0.65,1.80]$ and it has LSRK form again, but not its adjoint method. For the adjoint method, we now have in condition (33) for super-convergence [19, Lemma 5.1],
@@
if $\sigma_{n}\in[\underline{\sigma},\bar{\sigma}]$ and (34) are obeyed in the grid construction. The constant in (55) is computetd from $\hat{a}_{41}(1+\bar{\sigma}+\bar{\sigma}^{2})\doteq 0.611\leq 2/3$. This means that the adjoint method possesses global order for smooth grids only.
@@ 6.3 Global error estimates
Estimates for the global errors may be obtained by repeating the proof presented in [18] for uniform grids now for variable stepsizes using the weighted norm (25) constructed in the design process of the Peer triplets in [19]. Since its weight $W$ is the same on the whole grid with fixed condition number $\|W\|_{\infty}\|W^{-1}\|_{\infty}$, the estimate carries over to the maximum norm of the global error. By a fixed point argument, the proof in [18] showed that the maximal global error on the grid is bounded by the local errors in the form $3\|{\mathbb{M}}_{0}^{-1}\tau\|$, where the matrix ${\mathbb{M}}_{0}$ is a block matrix essentially containing the $h$-independent parts of the scheme. So, ${\mathbb{M}}_{0}^{-1}$ mainly consists of products of the stability matrices $\bar{B}_{n}=A_{n}^{-1}B_{n}$ or $A_{n}^{\sf T}B_{n+1}^{\sf T}$ which all are bounded by one in the norm (25) or its adjoint equivalent. Collecting variables as $z^{\sf T}=(y^{\sf T},p^{\sf T},u^{\sf T})$, $Z_{n}^{\sf T}=(Y_{N}^{\sf T},P_{n}^{\sf T},U_{n}^{\sf T})$, etc., and abbreviating $f(z):=f(y,u)$, the fixed-point argument requires upper limits on the derivatives $\nabla_{z}f,(1+\|z\|)\nabla_{zz}f$ in a tubular neighborhood of the exact solution $z^{\star}$. Moreover, a stronger version of the control uniqueness property mentioned above requires upper limits for $\big{\|}\big{(}p^{\sf T}\nabla_{uu}f(z^{\star})\big{)}^{-1}\big{\|}$ and $\|p^{\sf T}\nabla_{uzz}f(z)\|$ in the neighborhood.
Then, summing up the local errors in ${\mathbb{M}}_{0}^{-1}\tau$, the localized estimates from (28), (30) can be preserved. Since superconvergence eliminates the leading error not completely, the error bound will depend on two consecutive derivatives. The corresponding estimates will be abbreviated by a Matlab-type notation
$\|\varphi^{(q:r)}\|_{[n]}:=\max_{q\leq k\leq r}\|\varphi^{(k)}\|_{[n]},\quad 0\leq q\leq r,\ 0\leq n\leq N,$ (56)
for sufficiently smooth functions $\varphi(t)$. Combining now the proofs of Lemma 6.1 in [19] and Lemma 7.1 from [18], the global error may be estimated by
@@
$\leq\mu\max_{n=0}^{N}h_{n}^{s-1}\max\big{\{}\|(y^{\star})^{(s-1:s)}\|_{[n]},\|(p^{\star})^{(s-1:s)}\|_{[n]},\|(u^{\star})^{(s-1:s)}\|_{[n]}\big{\}},$ (57)
@@
where (57) holds for the pulcherrima triplet AP4o33vgi for general grids with bounded stepsize ratios $\sigma_{n}\in[\underline{\sigma},\bar{\sigma}]$ due to (52), while for AP4o33vsi also smoothness $|\sigma_{n}-1|\leq\eta h_{n}$ is required, see (55). We note that this error bound relates the local stepsize to the local size of solution derivatives as a justification for the kind of error control discussed in the next section.

### v2

@@ Abstract
This paper is concerned with the theory, construction and application of variable-stepsize implicit Peer two-step methods that are super-convergent for variable stepsizes, i.e., preserve their classical order achieved for uniform stepsizes when applied in a gradient-based solution algorithm to solve ODE-constrained optimal control problems in a first-discretize-then-optimize setting. Gradients of the objective function can be computed most efficiently using approximate adjoint variables. High accuracy with moderate computational effort can be achieved through time integration methods that satisfy a sufficiently large number of adjoint order conditions for variable stepsizes and provide gradients with higher-order consistency. In this paper, we enhance our previously developed variable implicit two-step Peer triplets constructed in [J. Comput. Appl. Math. 460, 2025] to get ready for large-scale dynamical systems with varying time scales without losing efficiency. A key advantage of Peer methods is their use of multiple stages with the same high stage order, which prevents order reduction — an issue commonly encountered in semi-discretized PDE problems with boundary control. Two third-order methods with four stages, good stability properties, small error constants, and a grid adaptation by equi-distributing global errors are constructed and tested for a 1D boundary heat control problem and an optimal control of cytotoxic therapies in the treatment of prostate cancer.
@@ 1 Introduction
The originality of this paper lies in the development of a full-fledged, practically deployable algorithm for large-scale ODE-constrained optimal control problems. To this end, the following modifications are implemented for two academic Peer triplets from [19]: (i) improved symmetry and stability properties and (ii) memory optimization in the start and end methods. At the same time, their classical third-order accuracy is preserved even for variable step sizes.
@@
The paper is organized as follows. In Section 2, we introduce a generalized form of a Peer two-step triplet. The continuous and discrete first order optimality conditions are presented in Section 3. The framework of gradient-based optimization, which relies on an iterative forward-backward marching scheme, is discussed in Section 4. Several properties of Peer triplets, including the standard and boundary methods, are described in Section 5. In Section 6, we modify two of our previous Peer two-step triplets of practical interest. Grid improvement is discussed in Section 7 and numerical results are given in Section 8. Conclusions are stated in Section 9.
Throughout the paper, we use the following notations: $e_{i}$ denotes a cardinal basis vector, $\mbox{1\thinspace l}_{k}=(1,\ldots,1)^{\sf T}\in{\mathbb{R}}^{k}$ is the vector of ones, and $I_{k}\in{\mathbb{R}}^{k\times k}$ is the identity matrix. The latter two are occasionally used without index if there is no ambiguity. Frequently used symbols are listed in Appendix B.
@@ 3 First order optimality conditions
Here, $p_{h}(T)=\nabla_{y}{{\cal C}}\big(y_{h}(T)\big)^{\sf T}$ and the Jacobians of $F$ are block diagonal matrices $\nabla_{Y}F(Y_{n},U_{n})=\,\mbox{blockdiag}_{i}\big(\nabla_{y}f(Y_{ni},U_{ni})\big)$. Further, the adjoint solutions $P_{n}=(P_{ni})_{i=1}^{s}\in{\mathbb{R}}^{sm}$ are approximations of $(p(t_{n}+c_{i}h_{n}))_{i=1}^{s}$. Since $h_{n}\kappa_{ii}>0$, (15) can be simplified to $\nabla_{u}f(Y_{ni},U_{ni})^{\sf T}P_{ni}\in N_{U}(U_{ni})$. Hence, the control uniqueness property mentioned above guarantees the existence of a local minimizer $U_{ni}$ of the Hamiltonian $H(Y_{ni},U,P_{ni})$ over all $U\in U_{ad}$, which shows the importance of having $\kappa_{ii}>0$ for all diagonal entries of $K$. Such positivity conditions also arise in the context of classical Runge-Kutta methods or W-methods, see [10, Theorem 2.1] and [20, Chapter 5.2].
@@ 5.1 The standard method
As for multi-step methods, order conditions are derived by Taylor expansion of the residuals if exact solution values are used as arguments of the numerical scheme. For variable stepsizes, these conditions depend on the stepsize ratio $\sigma_{n}=h_{n}/h_{n-1}$ in the scaling matrix $S_{n,q}=\,\mbox{diag}_{i}\big(\sigma_{n}^{i-1}\big)\in{\mathbb{R}}^{q\times q}$. Compact formulations of the order conditions are possible with the vector of node powers ${\bf c}^{k}=(c_{1}^{k},\ldots,c_{s}^{k})^{\sf T}$, the Vandermonde matrix $V_{q}=(\mbox{1\thinspace l},{\bf c},\ldots,{\bf c}^{q-1})\in{\mathbb{R}}^{s\times q}$, the Pascal matrix ${\cal P}_{q}=\big({j-1\choose i-1}\big)\in{\mathbb{R}}^{q\times q}$, and the scaled shift matrix $\tilde{E}_{q}=\big(i\delta_{i+1,j}\big)\in{\mathbb{R}}^{q\times q}$ which commutes with ${\cal P}_{q}=\exp(\tilde{E}_{q})$. Both the standard forward step (12) and its adjoint (14) posses local order $q$ if
@@
A critical requirement for the standard method is zero stability which may be very demanding for variable stepsizes. For the trivial ODE $y^{\prime}=0$, the stability matrix of the forward step (12), $Y_{n}=A^{-1}B_{n}Y_{n-1}$, of the standard method is $\bar{B}_{n}:=A^{-1}B_{n}$ and for the adjoint step (14) it is $\tilde{B}_{n+1}^{\sf T}:=A^{\sf-T}B_{n+1}^{\sf T}=A^{\sf-T}\bar{B}_{n+1}^{\sf T}A^{\sf T}$. Zero stability requires that arbitrarily long products, e.g., $\bar{B}_{n}\cdots\bar{B}_{k+1},\,n>k,$ are uniformly bounded. Since $\hat{A}^{-1}\hat{B}(\sigma)=V_{s}^{-1}\bar{B}(\sigma)V_{s}$ is similar to the stability matrix $\bar{B}(\sigma)$, based on the representation (22) in [19], the construction of a $\sigma$-independent weight matrix $W\in{\mathbb{R}}^{s\times s}$ was incorporated into the design of all standard methods, which block-diagonalizes all matrices simultaneously,
@@
$\tau_{n}^{Y}=$ $\;h_{n}^{q}\beta_{q,n}(\sigma_{n})(y^{\star})^{(q)}(t_{n})+O\left(h_{n}^{q+1}\|(y^{\star})^{(q+1)}\|_{[n]}\right),$ (28)
@@
$\tau_{n}^{P}=$ $\;h_{n}^{q}\beta_{q,n}^{\dagger}(\sigma_{n+1})(p^{\star})^{(q)}(t_{n})+O\left(h_{n}^{q+1}\|(p^{\star})^{(q+1)}\|_{[n]}\right),$ (30)
@@
see (23), with an analogous version for the adjoint matrices $\tilde{B}_{n+1}^{\sf T}$. Hence, there is no accumulation of leading error terms of the standard method under the super-convergence conditions
@@ 5.3 Triangular iterations at the boundaries
Remark 5.2
Replacing the matrix $A_{0}-h_{0}K\textbf{J}_{0}$ in the simplified Newton method by the lower block-triangular matrix $\tilde{A}_{0}-h_{0}K\textbf{J}_{0}$ with a small contraction factor for linear autonomous problems still delivers a fast converging sequence of approximate solutions. Thus, with appropriate stopping criteria, the overall third order of the Peer triplets is not affected by the triangular approximations at the boundaries.
@@ 6.2 The triplet AP4o33vsi for smooth grids
The diagonal matrix $K$ is also positive definite and, hence, preserves the sign in the cone condition (15). The triplet has a substantially larger angle $\alpha>83^{o}$ of $A(\alpha)$-stability than AP4o33vgi and the complement of its stability region is contained in $0.6,20]\times[-15,15]$. One drawback is larger error constants, see Table 1. The standard method is uniformly zero stable in $\underline{\sigma},\bar{\sigma}]=[0.65,1.80]$ with the modified weight matrix
$W=\begin{pmatrix}1&-\frac{7}{18}&-\frac{2}{21}&\frac{9}{14}\\[2.84526pt] 1&-\frac{47}{84}&\frac{11}{14}&-\frac{95}{126}\\[2.84526pt] 1&-\frac{3}{14}&-\frac{26}{63}&\frac{17}{30}\\[2.84526pt] 1&0&0&0\end{pmatrix}$.
The method has LSRK form again, but not its adjoint. For the adjoint method, we now have in condition (33) for super-convergence [19, Lemma 5.1],
@@
if $\sigma_{n}\in[\underline{\sigma},\bar{\sigma}]$ and (34) are obeyed in the grid construction. The constant in (55) is computetd from $\hat{a}_{41}(1+\bar{\sigma}+\bar{\sigma}^{2})\doteq 0.611\leq 2/3$. This means that the adjoint method possesses global order $q=3$ for smooth grids only.
@@ 6.3 Global error estimates
The global error in the unconstrained case $N_{U}=\{0\}$ may be derived by a slight generalization of the proof presented in [18] for uniform grids to the case of general grids obeying certain restrictions. The key point is that the design of the Peer methods includes the construction of fixed weight matrices $W,W^{\dagger}=(AW)^{\sf-T}$, where all stability matrices $\bar{B}_{n},\tilde{B}_{n+1}^{\sf T}$ are bounded by one with the possible exception of the two end steps, which do not spoil the estimates. Some smoothness of the problem is required, and with the abbreviations $z=(y^{\sf T},p^{\sf T},u^{\sf T})^{\sf T}$ and $f(z)=\big(f_{i}(z)\big)_{i=1}^{m}:=f(y,u)$, we assume bounds
$\|\nabla_{z}f(z)\|,(1+\|z\|)\|\nabla_{zz}f(z)\|\leq\Lambda_{2},\;\|\langle p,\nabla_{uzz}f(z)\rangle\|\leq\Lambda_{3},$ (56)
in a tubular neighborhood of the exact solution $z^{\star}(t)$. For the sake of clarity for derivatives including summations in ${\mathbb{R}}^{m}$, we introduced the notation $\langle p,\nabla_{z}f\rangle:=\sum_{i=1}^{m}(e_{i}^{\sf T}p)\nabla_{z}f_{i}$, etc. Since $K$ is a diagonal matrix in all steps now, the strengthened control uniqueness property [10] needs to be assumed at the exact solution only as
$\|\langle p^{\star}(t),\nabla_{uu}f(z^{\star}(t))\rangle^{-1}\|\leq\omega,\ t\in[0,T].$ (57)
The Peer methods exploit super-convergence, consequently, the global errors depend on more than one derivative of the solution. These estimates will be abbreviated by a Matlab-type notation
$\|\phi^{(q:r)}\|_{[n]}:=\max_{q\leq k\leq r}\|\phi^{(k)}\|_{[n]},\quad 0\leq q\leq r,\ 0\leq n\leq N,$ (58)
for sufficiently smooth functions $\phi(t)$. Only linear objective functions $\cal C$} are considered in order to shorten the proof.
Theorem 6.1
Let the solution $z^{\star}=\big((y^{\star})^{\sf T},(p^{\star})^{\sf T},(u^{\star})^{\sf T}\big)^{\sf T}$ of (6)–(8) be smooth with $y^{\star},p^{\star}\in C^{q+1}[0,T]$, let the objective function $\cal C$} in (1) be linear and $f$ satisfy the assumptions (56) in some tubular neighborhood of $z^{\star}$ and (57) in $z^{\star}$ itself. Assume that the Peer triplet satisfies the order conditions (18), (19) and (37), (39) with $q\geq 1$. Let there exist a fixed weight matrix $W\in{\mathbb{R}}^{s\times s}$ such that $\|W^{-1}A^{-1}B(\sigma)W\|_{\infty}=1$ and $\|W^{-1}A^{-1}B(\sigma)W-e_{1}e_{1}^{\sf T}\|_{\infty}\leq\tilde{\gamma}<1$ according to (23) for $\sigma\in[\underline{\sigma},\bar{\sigma}]$ with $0<\underline{\sigma}<1<\bar{\sigma}$. Then, under the condition
$\zeta_{1}T\Lambda_{2}\max\{1,4\omega\Lambda_{2}\}\leq\frac{2}{3},$ (59)
where $\zeta_{1}$ is a constant only depending on the Peer triplet, there exists a unique solution $Z=(Y^{\sf T},P^{\sf T},U^{\sf T})^{\sf T}$ to the system (11)–(15) for grids $\{t_{n}\}$ with stepsize ratios $\sigma_{n}\in[\underline{\sigma},\bar{\sigma}]$ and $H=\max_{n=0}^{N}h_{n}$ small enough. Under the additional condition (52) for super-convergence, the error to the exact solution ${\bf z}_{n}^{\sf T}=({\bf y}_{n}^{\sf T},{\bf p}_{n}^{\sf T},{\bf u}_{n}^{\sf T})$ is bounded as
@@
$\leq\mu\max_{n=0}^{N}h_{n}^{q}\max\big\{\|(y^{\star})^{(q:q+1)}\|_{[n]},\|(p^{\star})^{(q:q+1)}\|_{[n]}\big\}.$ (60)
with some constant $\mu$. The estimate (60) also holds under the weaker condition (55) for smooth grids obeying $\sigma_{n}=1+O(h_{n})$, see (34).
@@
We note that the Theorem applies to the Peer triplets AP4o33vgi, AP4o33vsi with q=s1=3q=s-1=3 and that the error bound relates the local stepsize to the local size of solution derivatives as a justification for the kind of error control discussed in the next section.
Sketch of proof similar to [18], Theorem 7.2: Errors are denoted as $\check{Z}_{n}=Z_{n}-{\bf z}_{n}$. Subtracting in a usual way the equations (12), (14) defining the standard method, and those for the truncation error, (26), (27), we get for the standard method the error equations
$\check{Y}_{n}-\bar{B}_{n}\check{Y}_{n-1}$ $=R_{n}^{Y}(\check{Z})-\tau_{n}^{Y},\;n=1,\ldots,N,$ (61)
$R_{n}(\check{Z})$ $:=h_{n}\bar{K}_{n}\big(F({\bf z}_{n}+\check{Z}_{n})-F({\bf z}_{n})\big),$
$\check{P}_{n}-\tilde{B}_{n+1}^{\sf T}\check{P}_{n+1}$ $=R_{n}^{P}(\check{Z}_{n})-\tau_{n}^{P},\;n=0,\ldots,N-1,$ (62)
$R_{n}^{P}(\check{Z})$ $:=h_{n}A_{n}^{\sf-T}\big(\langle K({\bf p}_{n}+\check{P}_{n}),\nabla_{Y}F({\bf z}_{n}+\check{Z}_{n})\rangle-\langle K{\bf p}_{n},\nabla_{Y}F({\bf z}_{n}\rangle\big),$
with $\bar{B}_{n}=A_{n}^{-1}B_{n}$, $\bar{K}_{n}=A_{n}^{-1}K$, $\tilde{B}_{n+1}^{\sf T}=A_{n}^{\sf-T}B_{n+1}^{\sf T}$. In the equations (61), (62), the linear parts with constant coefficients were collected on the left hand sides. The error equations for the starting step (11) and the adjoint boundary condition (13) are treated in a similar manner. The constraint (15) with $N_{U}=\{0\}$ reads $\kappa_{ii}\langle P_{ni},\nabla_{u}f(Z_{ni})\rangle=0$, $n=0,\ldots,N$, $i=1,\ldots,s$, and has exactly the same form as (8), which means that there is no truncation error here, i.e., $\tau_{n}^{U}=0$, and no smoothness assumptions on the control are required. In order to obtain the same structure for this constraint with constant coefficients on the left hand side as in (61), (62), we have to fix derivatives at the exact solution ${\bf z}_{n}$. Also $\kappa_{ii}>0$ may be dropped in these conditions. So for the total global error $\check{Z}=\big(\check{Z}_{n}\big)_{n=0}^{N}$, we end up with the system
${\mathbb{M}}\check{Z}=-\tau+R(\check{Z}):=\begin{pmatrix}-\tau^{Y}+R^{Y}(\check{Z})\\ -\tau^{P}+R^{P}(\check{Z})\\ R^{U}(\check{Z})\end{pmatrix},\quad{\mathbb{M}}=\begin{pmatrix}M_{11}\otimes I_{m}&0&0\\ 0&M_{22}\otimes I_{m}&0\\ {\mathbb{M}}_{31}&{\mathbb{M}}_{32}&\Omega\end{pmatrix}$, (63)
where ${\mathbb{M}}_{31}=\mbox{blockdiag}_{n,i}\big(\langle{\bf p}_{ni},\nabla_{uy}f({\bf z}_{ni})\rangle\big)$, ${\mathbb{M}}_{32}=\mbox{blockdiag}_{n,i}\big(\nabla_{u}F({\bf z}_{ni})^{\sf T}\big)$, and in the main diagonal, $\Omega=\mbox{blockdiag}_{n,i}\big(\langle{\bf p}_{ni},\nabla_{uu}f({\bf z}_{ni})\rangle\big)$. Protracted computations in [18] showed Lipschitz conditions for the right-hand sides $R_{n}$ of the form
$\|R_{n}^{Y}(\tilde{Z})-R_{n}^{Y}(\hat{Z})\|,\,\|R_{n}^{P}(\tilde{Z})-R_{n}^{P}(\hat{Z})\|\leq h_{n}\zeta_{1}\Lambda_{2}\|\tilde{Z}_{n}-\hat{Z}_{n}\|,$ (64)
$\|R_{n}^{U}(\tilde{Z})-R_{n}^{U}(\hat{Z})\|\leq(\zeta_{2}\Lambda_{2}+\zeta_{3}\Lambda_{3})(\|\tilde{Z}_{n}\|+\|\hat{Z}_{n}\|)\|\tilde{Z}_{n}-\hat{Z}_{n}\|,$ (65)
for $\tilde{Z}_{n},\hat{Z}_{n}$ in neighborhoods of the origin. We note that the constants $\zeta_{i}$ only depend on the coefficients of the Peer method. The matrices $M_{11},M_{22}$ in (63) are block bidiagonal matrices with identities $I_{s}$ in the main diagonal. According to (61), $M_{11}$ has $\bar{B}_{n}$ in its first subdiagonal blocks while the first superdiagonal blocks in $M_{22}$ are $\tilde{B}_{n+1}^{\sf T}$ by (62). We recall that both satisfy $\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\bar{B}_{n}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert=\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\tilde{B}_{n+1}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert=1$ in the respective weighted norms. Hence, rewriting (63) in fixed point form $\check{Z}=\Phi(\check{Z})$ with $\Phi(\check{Z})={\mathbb{M}}^{-1}(R(\check{Z})-\tau)$, we obtain in these weighted norms for the first two parts $\Phi^{Y},\Phi^{P}$ the estimates
$\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\Phi^{Y}(\tilde{Z})-\Phi^{Y}(\hat{Z})\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert,\,\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\Phi^{P}(\tilde{Z})-\Phi^{P}(\hat{Z})\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\leq T\zeta_{1}\Lambda_{2}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\tilde{Z}-\hat{Z}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert$, (66)
where $\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert Z\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert=\max_{n=0}^{N}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert Z_{n}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert$. Now, the Lipschitz condition for the last row of (63) may be rewritten as
[…]
leading with assumption (57) and (66) to
$\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\Phi^{U}(\tilde{Z})-\Phi^{U}(\hat{Z})\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\leq 2\omega((\zeta_{2}\Lambda_{2}+\zeta_{3}\Lambda_{3})\varepsilon+\zeta_{1}T\Lambda_{2}^{2}\big)\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\tilde{Z}-\hat{Z}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert$
[…]
$\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\check{Z}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert=3\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\Phi(\check{Z})\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert-2\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\check{Z}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\leq 3\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\Phi(\check{Z})-\Phi(0)\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert-2\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\check{Z}\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert+3\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\Phi(0)\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\leq 3\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert{\mathbb{M}}^{-1}\tau\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert$ (67)
under the assumption $3\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert{\mathbb{M}}^{-1}\tau\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert\leq\hat{\varepsilon}$.
Applying simple summation of bounds in $\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert{\mathbb{M}}^{-1}\tau\rvert\hskip-1.19995pt\rvert\hskip-1.19995pt\rvert$ would lose one order of convergence. This loss can be avoided with the aid of the super-convergence conditions (52), (55) along the lines of the proof of Lemma 7.3 in [18]. According to (23), long products $\bar{B}_{n}\cdots\bar{B}_{k-1},\,k<n,$ converge to the rank one matrix $\mbox{1\thinspace l}\mbox{1\thinspace l}^{\sf T}A$ as $\|W^{-1}\bar{B}_{n}\cdots\bar{B}_{k+1}W-e_{1}e_{1}^{\sf T}\|_{\infty}\leq\tilde{\gamma}^{n-k}$. Now, due to $\mbox{1\thinspace l}^{\sf T}A\beta_{3}(\sigma)=e_{1}^{\sf T}W^{-1}\beta_{3}(\sigma)\equiv 0$, the contribution of, e.g., $M_{11}^{-1}\tau^{Y}$ in (67) for time step $n$ is
$\|W^{-1}\sum_{k=0}^{n}\bar{B}_{n}\cdots\bar{B}_{k+1}\tau_{k}^{Y}\|_{\infty}\leq\|\sum_{k=0}^{n}e_{1}e_{1}^{\sf T}W^{-1}\tau_{k}^{Y}\|_{\infty}+\sum_{k=0}^{n}\tilde{\gamma}^{n-k}\|\tau_{k}^{Y}\|_{\infty}$
$=O\big(\sum_{k=0}^{n}(h_{k}^{q+1}\|(y^{\star})^{(q+1)}\|_{[n]}+\tilde{\gamma}^{n-k}h_{k}^{q}\|(y^{\star})^{(q)}\|_{[n]})\big)=O\big(\max_{k=0}^{n}h_{k}^{q}\|(y^{\star})^{(q:q+1)}\|_{[k]}\big).$ (68)
Hence, for $q\geq 1$ and $H$ small enough, the estimate (68) will be smaller than $\hat{\varepsilon}/3$ and (67) yields the assertion. For the adjoint error $M_{22}^{-1}\tau^{P}$ of the Peer method AP4o33vsi, where $\beta_{3}^{\dagger}$ satisfies (55) only, the leading error in $e_{1}^{\sf T}(W^{\dagger})^{-1}\tau_{k}^{P}$ is now $O(h_{k}^{q+1}\|(p^{\star})^{(q:q+1)}\|_{[n]})$ for smooth grids, which is covered by the final bound (60). Here, $W^{\dagger}=(AW)^{\sf-T}$ is the weight matrix for all matrices $\tilde{B}_{n+1}^{\sf T}$.   $\square$